OpenStudy (anonymous):

Kendrick bought 2 identical pizzas . He cut the first pizza into 5 equal pieces and the second pizza in to 3 equal pieces. He ate 2 pieces from each pizza what was the total mass of the pizzas eaten if the mass of the remaining pizzas was 540grams ?

OpenStudy (ddcamp):

What fraction of the first pizza was left? The second? Add those together. Let that be the variable l. l × (total mass) =540g

OpenStudy (anonymous):

I do not get you

OpenStudy (ddcamp):

So, he ate 2/5 pieces of the first pizza, right? And 2/3 of the second: (2/5) + (2/3) = (6/15) + (10/15) = (16/15) Then, divide by 2 (because there were two pizzas): 8/15 overall of the pizza. So, 7/15 were left. If 7/15 of the total = 540g, what would 8/15 be?

OpenStudy (anonymous):

I am very sorry error in question it is 504grams not 540grams but is still the method the same ?